# Visual Semantic Navigation with Real Robots

### v1

@@ 3.1.4 Visual Semantic Navigation package
This submodule is in charge of collecting odometry information by subscribing to the /odom topic (odometry topic), published by the robot_api package. This odometry information consists of two main variables: 1) robot position, that indicates the current location of the robot with respect to its initial position; and 2) robot orientation, that defines the current direction in which the robot is oriented in relation to its initial orientation. Some state-of-the-art VSN models (e.g. (Ramrakhya et al., 2023)) need to input these two sources of information.
@@
A configuration file is provided (vsn.yaml) containing default values for the parameters of our VSN package. By modifying this file one can easily change the navigation target, parameters associated with the median filter, or even the maximum number of steps allowed to be executed during a navigation exercise.
@@ 3.2 VSN models
For this research work, we have decided to adapt and integrate into our robots two state-of-the-art VSN models: PIRLNav (Ramrakhya et al., 2023) and VLV (Chang et al., 2020). The first model, known as PIRLNav (Ramrakhya et al., 2023), is a VSN approach that has been trained with a combination of imitation learning and a RL fine-tuning. As of today, this model reports the best results in the ObjectNav (Batra et al., 2020) task in Habitat (Szot et al., 2021). The second model is the VLV approach (Chang et al., 2020), which is a VSN model directly trained from YouTube videos. The VLV model makes use of such videos to learn semantic cues for an effective navigation to semantic targets in indoor scenarios. VLV is a modular learning solution, that combines low-level and high-level navigation policies.
@@ 3.2.1 VLV
The first approach is known as Value Learning from Videos (VLV), developed by Chang et al. (2020). VLV is a modular-learning based VSN model directly trained from videos of real state agencies, taken from YouTube. In this type of video, a human records, camera in hand, the properties for sale, showcasing all the rooms they have, to generate a sort of virtual tour of the houses. Note that the videos used do not have any type of information about the navigation actions that take place during the recording, nor on what kind of rooms or what type of objects appear.
@@
The VLV model leverages such YouTube videos to learn semantic cues for an effective navigation to semantic targets in indoor home environments. This way, the VLV model is trained to find in these videos a set of object categories. Technically, the model uses pseudo action labels obtained by running an inverse model on the navigation sequences. This inverse model is able to recognize the discrete movements that each of the transitions of the different video frames involve. Then, the navigation policies are learned following a reinforcement learning (RL) approach. VLV employs Q-learning to learn from the video sequences that have been pseudo-labeled with the actions. The learned Q-function, and the associated value function, implicitly learn semantic cues for navigation. In other words, the model learns what images lead to the desired category, and what do not.
@@
For our experiments, we had to embed the VLV model in our ROS4VSN architecture. See Figure 4(a). Technically, we integrated the two navigation policies detailed in the experiments in (Chang et al., 2020) that were tested in the virtual environment Habitat (Szot et al., 2021). However, now, our objective is to implement them in a real robot in the real world.
@@
This integration into our robots, using our ROS4VSN architecture, has consisted of the following steps. First, a high-level policy that stores 12 images for each node in a topological graph (obtained by rotating 12 times by 30 degrees each) is employed. This high-level policy uses the learned value function score over these 12 images, and samples the most promising direction for seeking objects of a particular object category. VLV needs an object detector output to produce the final score for these images. This way, the high-level policy is equipped with a mechanism to seek the object once it has been detected. Specifically, the detector we employ is the Mask R-CNN (He et al., 2017), which we had to embed in our architecture as well. To navigate, this policy will use the following discrete movements/actions: move_forward 25cm, turn_right $30^{\circ}$ or turn_left $30^{\circ}$.
@@
These movements are compatible with the developed discrete_move package of our ROS4VSN architecture. Once a main direction has been chosen, our approach converts it into a short-term goal by sampling a location at an offset of 1.5 meters from the chosen node’s location, in the chosen view’s direction. This is done using the depth-camera and a low-level navigation policy that uses occupancy maps with a fast marching planning (Sethian, 1996) to execute robot actions to reach the short-term goal. These two policies have been integrated into our ROS4VSN nodes, and with them we get the robot to explore the environment.
@@ 3.2.2 PIRLNav
The second model selected for our experimental evaluation is known as PIRLNav (Ramrakhya et al., 2023). As of today, this model reports the best performance in the ObjectNav (Batra et al., 2020) task in Habitat (Szot et al., 2021).
Figure 6: One of the 3D reconstructions of the HM3D-Semantics v0.1 dataset (Ramakrishnan et al., 2021), used to train PILRNav.
@@
PILRNav is a VSN approach that has been trained with a combination of imitation learning and a RL fine-tuning. The model uses behavior cloning (BC) to pre-train the ObjectNav policy on a dataset of $77k$ human demonstrations, amounting more than $2370$ human annotation hours, in the HM3D-Semantics v0.1 dataset (Ramakrishnan et al., 2021). This dataset provides up to 120 different 3D reconstructions of houses all around the world (see Figure 6 for an example of one of them). Once this BC is finished, a RL finetuning is used following the DD-PPO approach (Wijmans et al., 2019). The policy architecture used is a simple CNN plus a RNN from (Yadav et al., 2023b).
@@
In order to integrate PIRLNav into our ROS4VSN architecture, we had to perform the following actions. See figure 4(b). The original PIRLNav needs to receive as inputs the RGB image that the agent observes, as well as the noiseless GPS and compass information offered by Habitat simulator. GPS and compass Habitat sensor provides the agent’s current location and orientation information relative to the start of the episode. In our case, because PIRLNav has to be integrated in a real robot, navigating in the real world, we proceed to feed the model with the RGB images that are acquired by the cameras in our robotic platforms. GPS information is obtained through the odometry information provided by the robot. For the compass, we recover the relative orientation analyzing all the robot’s turning movements. Note that these are not anymore noiseless sensors, as the ones used in the simulated world in which the PIRLNav model was trained. Fortunately, we did not observe any important impact on the performance of the model, due to this loss of precision for these sensors.
@@
PIRLNav is therefore integrated in the ROS4VSN architecture to control de navigation of the robot as it has been detailed. For every captured image, as well as the GPS+compass data, the model is able to determine the next discrete movement action to be executed by our robotic platforms. The action space used for our experiments with this model is: move_forward 25 cm, move_backward 25 cm, turn_right $30^{\circ}$, turn_left $30^{\circ}$ and stop. The original PIRLNav was trained to produce also the discrete action look_up and look_down, since the simulated agent could tilt its camera. However, as it is observed in Figure 8, in our platforms these actions are not possible. We decided to replace look_up with a move_backward action, and look_down movement with the move_forward action. This choice is based on the reasoning that, by raising the camera, more of the scene is captured; moving the robot backward serves this purpose. Also, since lowering the camera provides a greater level of scene detail, moving forward is considered the most appropriate choice to replace the look_down action. Finally, in order to prevent collisions between the robot and objects in the scene, a procedure was developed that uses information from the depth image to detect obstacles at a given distance. Note that PIRLNav does not need any low-level policy as in the VLV model. The robot is controlled and navigates using only the set of discrete actions provided by the ROS4VSN model.
@@ 4.1 Experimental setup
One of the main objectives of our work is to provide a thorough evaluation of state-of-the-art VSN models in a real scenario with real robots. For doing so, we propose the following experimental setup.
@@
[…]
From these positions, the robot is asked to navigate towards the different object categories. Therefore, we perform 15 navigation experiments for every target and measure the success of the episodes if the robot reaches the selected object category in less than 150 discrete actions, and without collisions. This limit of 150 steps was chosen to establish a balance between the average size of the houses typically used in Habitat (Szot et al., 2021) and the apartment used in our experiments. We report the success rate (SR) of the VSN models as the percentage of episodes in which navigation was successful. An experiment is considered successful if the robot stops (the VSN model samples the action stop) and the Euclidean distance to the target object is less than one meter. Note that our navigation experiment mimics the evaluation performed in the ObjectNav (Batra et al., 2020) task, with the same metric. This is the standard experiment on which most VSN models are compared and which currently defines the state of the art.
We have used two different robots for our experiments: a Turtlebot 2, and the LOLA2 (Nasri et al., 2022) platform. For doing so, we had to embed our ROS4VSN in both of these platforms. This can be easily done adapting the robot_api module described in Section 3.1.1. Figure 8 shows the family picture of the robots invited to our experiments. We mainly used the Turtlebot 2 for the navigation experiments in the apartment described. The robot LOLA2 was also used in navigation experiments, but in a different location, in order to test the stability of the developed system, and to provide a study that contains more hours of navigation, and on different platforms. Our intention in using two different platforms has been to provide evidence of the generalizability of our architecture, demonstrating that it can be tested on different robots.
@@
When collecting information during the experiments, we developed a procedure to record relevant data for each trial. This procedure stores the unique identifier for each episode, the sequence of actions performed and the category of object searched. In addition, during the tests, qualitative information about the trajectory followed by the robot was recorded. In particular, all the images observed by the robot during its trajectories have been saved.
@@
For the experiments, the following hardware-software setup has been employed, as it is shown in Figure 9. The modular architecture of the developed ROS4VSN system was used to deploy it in a distributed manner. This architecture allows separating the execution of ROS packages on different devices, as long as they are connected to the same network. The robotic platforms were equipped with a laptop. This device was used to establish the communication with the robotic platform and the camera by executing the robot_api and camera_api packages. At the same time, the discrete_move package was executed to receive the actions to be executed by the robot. On the other hand, we used a workstation to run the VSN nodes, which was equipped with an i7-1165G7 processor and an NVIDIA GeForce RTX 2060 graphics card.
@@ 4.2 VSN navigation results
We provide in this section the main results obtained during the navigation experiments with our robots, including both quantitative and qualitative results.
Following the experimental setup detailed in Section 4.1, we have obtained the following results. Remember that we provide in this study an analysis of the SR for two state-of-the-art VSN models, running in a Turtlebot 2 platform. For every VSN model, i.e. VLV and PIRLNav, we have measured their SR for the different object categories they can navigate to.
Table 2 compares the performance obtained by VLV and PIRLNav approaches when they are tested in the real world (i.e. our experiments) and in a virtual environment (i.e. the experiments reported in their respective papers). The first thing we observe is the difference in terms of SR. The SR for the PIRLNav model drops from $65\%$ to $21\%$, while the VLV model loses $\sim 10$ percentage points in this metric. One of the conclusions of our study is that there is a considerable gap between the behavior of these models in the real world and in the simulation environments in which they are trained. This indicates that further research in this direction is needed. Interestingly, the results obtained in our real-world experiments are not consistent with the performance difference that already existed between the models in the simulation environments: VLV is the winner in the real world! As we analyze in the discussion section below (See section 4.4), we believe that this behavior is due to the impact of the object detector that VLV integrates, but PIRLNav does not. While the difference between VLV and PIRLNav SR in the virtual environments is of $26$ percentage points, in the real world this gap becomes of just $8$ percentage points.
@@
In the following, we analyze in detail the results reported by each of the models. We start with VLV model. Table 3 reports the SR for every of the target categories used in our experiments. Chairs, tables and sofas are the categories that are easiest to navigate to. In analyzing various trials with the robot using the VLV model with ROS4VSN, there were no successful outcomes from starting positions 10 and 12 (see Figure 7). Additionally, only one success was observed from positions 1, 2, and 13. Notably, our VLV implementation can reach most targets in under 60 steps.
@@
Table 3: VLV VSN experiment. We report the number or successful episodes over 15, the corresponding SR per object goal, and the average number of actions taken to reach the target.
Figure 10 shows qualitative results for four navigation experiments with VLV model. We provide five representative images of the navigation experiments. Two successful and two unsuccessful cases are presented. In the first experiment (first row), the robot quickly reached the Table, as the detector easily identified it in the images. In the second experiment, the robot took a detour to the Chair because the detector failed to detect it initially. The model predicted a point near the chair but out of view, causing the robot to move closer only after it became visible. In the third and fourth experiments, the robot started in a challenging position in front of a refrigerator. Due to noise in the depth image, the target point calculation was inaccurate, leading to collisions with the wall in both cases.
@@
We analyze now in detail the results reported by PIRLNav model. Table 4 shows the SR obtained for the PIRLNav model integrated into our ROS4VSN architecture. We detail the SR reported for every object category. The agent was able to more easily locate the most common objects in the house, such as the chair and the monitor. The abundant and well-distributed presence of these objects facilitated the agent’s work. The model inferred the action stop on both objects a total of five times. Large objects, such as the sofa and bed, showed slightly lower results. Although easily visible from multiple locations in the dwelling, the presence of only one of these objects made it difficult for the robotic agent to spot them. The number of times the stop action was sampled for these categories was substantially reduced. With the toilet we have one of the most complex challenges in navigating the robot in this home. The robot did not have full visibility of the toilet until it managed to fully enter the bathroom, having to pass through the narrow door without colliding. Finally, for the category plant, the robot was not able to locate this category in any of the 15 attempts. Even though the plant was visible on multiple occasions during navigation, the agent did not manage to head towards this object. Overall, considering all the categories, the SR for the PIRLNav model is of $21.11\%$.
@@
Table 4: PIRLNav VSN experiment. We report the number or successful episodes over 15, the corresponding SR per object goal, and the average number of actions taken to reach the target.

### v2

@@ 3.1.4 Visual Semantic Navigation package
This submodule is in charge of collecting odometry information by subscribing to the /odom topic (odometry topic), published by the robot_api package. This odometry information consists of two main variables: 1) robot position, that indicates the current location of the robot with respect to its initial position; and 2) robot orientation, that defines the current direction in which the robot is oriented in relation to its initial orientation. Some state-of-the-art VSN models (e.g [1]) need to input these two sources of information.
@@
A configuration file is provided (vsn.yaml) containing default values for the parameters of our VSN package. By modifying this file, one can easily change the navigation target, the parameters associated with the median filter, or even the maximum number of steps allowed to be executed during a navigation exercise.
@@ 3.2 VSN models
For this research work, we have decided to adapt and integrate into our robots two state-of-the-art VSN models: PIRLNav [1] and VLV [3]. The first model, known as PIRLNav [1], is a VSN approach that has been trained with a combination of imitation learning and a RL fine-tuning. As of today, this model reports the best results in the ObjectNav [4] task in Habitat [55]. The second model is the VLV approach [3], which is a VSN model directly trained from YouTube videos. The VLV model makes use of such videos to learn semantic cues for an effective navigation to semantic targets in indoor scenarios. VLV is a modular learning solution that combines low-level and high-level navigation policies.
@@ 3.2.1 VLV
The first approach is known as Value Learning from Videos (VLV), developed by Chang et al. [3]. VLV is a modular-learning based VSN model directly trained from videos of real state agencies, taken from YouTube. In this type of video, a human records, camera in hand, the properties for sale, showcasing all the rooms they have, to generate a sort of virtual tour of the houses. Note that the videos used do not have any type of information about the navigation actions that take place during the recording, nor in what kind of rooms or what type of objects appear.
@@
The VLV model leverages such YouTube videos to learn semantic cues for an effective navigation to semantic targets in indoor home environments. This way, the VLV model is trained to find in these videos a set of object categories. Technically, the model uses pseudo action labels obtained by running an inverse model on the navigation sequences. This inverse model is able to recognize the discrete movements that each of the transitions of the different video frames involve. Then, the navigation policies are learned following a reinforcement learning (RL) approach. VLV employs Q-learning to learn from the video sequences that have been pseudo-labeled with the actions. The learned Q-function, and the associated value function, implicitly learn semantic cues for navigation. In other words, the model learns what images lead to the desired category and what do not.
@@
For our experiments, we had to embed the VLV model in our ROS4VSN architecture. See Figure 5(a). Technically, we integrated the two navigation policies detailed in the experiments in [3] that were tested in the virtual environment Habitat [55]. However, now, our goal is to implement them in a real robot in the real world.
@@
This integration into our robots, using our ROS4VSN architecture, has consisted of the following steps. First, a high-level policy that stores 12 images for each node in a topological graph (obtained by rotating 12 times by 30 degrees each) is used. This high-level policy uses the learned value function score over these 12 images, and samples the most promising direction for seeking objects of a particular object category. VLV needs an object detector output to produce the final score for these images. This way, the high-level policy is equipped with a mechanism to seek the object once it has been detected. Specifically, the detector we employ is the Mask R-CNN [56], which we had to embed in our architecture as well. To navigate, this policy will use the following discrete movements/actions: move_forward 25cm, turn_right $30^{\circ}$ or turn_left $30^{\circ}$.
@@
These movements are compatible with the developed discrete_move package of our ROS4VSN architecture. Once a main direction has been chosen, our approach converts it into a short-term goal by sampling a location at an offset of 1.5 meters from the chosen node's location, in the chosen view's direction. This is done using the depth-camera and a low-level navigation policy that uses occupancy maps with a fast marching planning [57] to execute robot actions to reach the short-term goal. These two policies have been integrated into our ROS4VSN nodes, and with them, we get the robot to explore the environment.
@@ 3.2.2 PIRLNav
The second model selected for our experimental evaluation is known as PIRLNav [1]. As of today, this model reports the best performance in the ObjectNav [4] task in Habitat [55].
Figure 6: One of the 3D reconstructions of the HM3D-Semantics v0.1 dataset [58], used to train PILRNav.
@@
PILRNav is a VSN approach that has been trained with a combination of imitation learning and a RL fine-tuning. The model uses behavior cloning (BC) to pre-train the ObjectNav policy on a dataset of $77k$ human demonstrations, amounting more than $2370$ human annotation hours, in the HM3D-Semantics v0.1 dataset [58]. This dataset provides up to 120 different 3D reconstructions of houses all around the world (see Figure 6 for an example of one of them). Once this BC is finished, a RL fine-tuning is used following the DD-PPO approach [59]. The policy architecture used is a simple CNN plus an RNN from [17].
@@
In order to integrate PIRLNav into our ROS4VSN architecture, we had to perform the following actions. See figure 5(b). The original PIRLNav needs to receive as inputs the RGB image that the agent observes, as well as the noiseless GPS and compass information offered by Habitat simulator. GPS and compass Habitat sensor provide the agent's current location and orientation information relative to the start of the episode. In our case, because PIRLNav has to be integrated in a real robot, navigating in the real world, we proceed to feed the model with the RGB images that are acquired by the cameras in our robotic platforms. GPS information is obtained through the odometry information provided by the robot. For the compass, we recover the relative orientation analyzing all the robot’s turning movements. Note that these are not anymore noiseless sensors, as the ones used in the simulated world in which the PIRLNav model was trained. Fortunately, we did not observe any important impact on the performance of the model, due to this loss of precision for these sensors.
@@
PIRLNav is therefore integrated in the ROS4VSN architecture to control de navigation of the robot as it has been detailed. For every captured image, as well as the GPS+compass data, the model is able to determine the next discrete movement action to be executed by our robotic platforms. The action space used for our experiments with this model is: move_forward 25 cm, move_backward 25 cm, turn_right $30^{\circ}$, turn_left $30^{\circ}$ and stop. The original PIRLNav was also trained to produce the discrete action look_up and look_down, since the simulated agent could tilt its camera. However, as it is observed in Figure 8, in our platforms these actions are not possible. We decided to replace look_up with a move_backward action, and look_down movement with the move_forward action. This choice is based on the reasoning that, by raising the camera, more of the scene is captured; moving the robot backward serves this purpose. Also, since lowering the camera provides a greater level of scene detail, moving forward is considered the most appropriate choice to replace the look_down action. Finally, to prevent collisions between the robot and objects in the scene, a procedure was developed that uses information from the depth image to detect obstacles at a given distance. Note that PIRLNav does not need any low-level policy as in the VLV model. The robot is controlled and navigates using only the set of discrete actions provided by the ROS4VSN model.
@@ 4.1 Experimental setup
One of the primary objectives of our work is to provide a comprehensive and clear protocol for the experimental evaluation of state-of-the-art VSN models in real-world scenarios using real robots. Our goal is for other researchers to carry out similar experimental evaluations, using the same evaluation metrics, to facilitate comparisons of how different VSN models perform in real-world navigation tasks. For doing so, we propose the following experimental setup.
@@
In a 75 m2 apartment, we define up to 15 random starting positions (see Fig. 7). The apartment can be divided into three main areas: a bedroom, a bathroom, and a larger space that includes the kitchen, living room, and study area. This setting contains all the object categories used to train the VSN models in the experiments, such as chair, bed, plant, bathroom, monitor, table, and sofa. We encourage other researchers to conduct their experiments in real-world settings that are similar in size and characteristics, allowing the robot to navigate from at least 15 different starting positions across multiple instances.
From these positions, the robot is tasked with navigating to various object categories. Consequently, one must conduct 15 navigation experiments for each target category and measure the success of the episodes based on whether the robot reaches the designated object category in fewer than 150 discrete actions and without any collisions. This limit of 150 steps was chosen to establish a balance between the average size of the houses typically used in Habitat [55] and the apartment used in our experiments.
For the evaluation metric, we propose reporting the success rate (SR) of the VSN models as the percentage of episodes in which navigation is deemed successful. An experiment is considered successful if the robot halts (when the VSN model samples the action stop) and the Euclidean distance to the target object is less than one meter.
Note that our navigation experiment mimics the evaluation performed in the ObjectNav [4] task, with the same metric. This is the standard experiment on which most VSN models are compared and which currently defines the state of the art.
We have used two different robots for our experiments: a Turtlebot 2, and the LOLA2 [53] platform. To do so, we had to embed our ROS4VSN in both of these platforms. This can be easily done by adapting the robot_api module described in Section 3.1.1. Figure 8 shows the family picture of the robots invited to our experiments. We mainly used the Turtlebot 2 for the navigation experiments in the apartment described. The robot LOLA2 was also used in navigation experiments, but in a different location, to test the stability of the developed system and to provide a study that contains more hours of navigation, and on different platforms. Our intention in using two different platforms has been to provide evidence of the generalizability of our architecture, demonstrating that it can be tested on different robots.
@@
When collecting information during the experiments, we developed a procedure to record relevant data for each trial. This procedure stores the unique identifier for each episode, the sequence of actions performed, and the category of object searched. In addition, during the tests, qualitative information about the trajectory followed by the robot was recorded. In particular, all the images observed by the robot during its trajectories have been saved.
@@
For the experiments, the following hardware-software setup has been used, as it is shown in Figure 9. The modular architecture of the developed ROS4VSN system was used to deploy it in a distributed manner. This architecture allows separating the execution of ROS packages on different devices, as long as they are connected to the same network. The robotic platforms were equipped with a laptop. This device was used to establish the communication with the robotic platform and the camera by executing the robot_api and camera_api packages. At the same time, the discrete_move package was executed to receive the actions to be executed by the robot. On the other hand, we used a workstation to run the VSN nodes, which was equipped with an i7-1165G7 processor and an NVIDIA GeForce RTX 2060 graphics card. We provide code to reproduce all our experiments at https://github.com/gramuah/ros4vsn.
@@ 4.2 VSN navigation results
We detail in this section the main results obtained during the navigation experiments with our robots, including both quantitative and qualitative results.
Following the experimental setup detailed in Section 4.1, we have obtained the following results. Remember that we provide in this study an analysis of the SR for two state-of-the-art VSN models, running in a Turtlebot 2 platform. For every VSN model, i.e VLV, and PIRLNav, we have measured their SR for the different object categories they can navigate to.
Table 2 compares the performance obtained by VLV and PIRLNav approaches when they are tested in the real world (i.e our experiments) and in a virtual environment (i.e the experiments reported in their respective papers). The first thing we observe is the difference in terms of SR. The SR for the PIRLNav model drops from $65\%$ to $21\%$, while the VLV model loses $\sim 10$ percentage points in this metric. One of the conclusions of our study is that there is a considerable gap between the behavior of these models in the real world and in the simulation environments in which they are trained. This indicates that further research in this direction is needed. Interestingly, the results obtained in our real-world experiments are not consistent with the performance difference that already existed between the models in the simulation environments: VLV is the winner in the real world! As we analyze in the discussion section below (See section 4.4), we believe that this behavior is due to the impact of the object detector that VLV integrates, but PIRLNav does not. While the difference between VLV and PIRLNav SR in the virtual environments is of $26$ percentage points, in the real world this gap becomes of just $8$ percentage points.
@@
In the following, we analyze in detail the results reported by each of the models. We start with the VLV model. Table 3 reports the SR for every of the target categories used in our experiments. Chairs, tables, and sofas are the categories that are easiest to navigate to. In analyzing various trials with the robot using the VLV model with ROS4VSN, there were no successful outcomes from starting positions 10 and 12 (see Figure 7). Additionally, only one success was observed from positions 1, 2, and 13. Notably, our VLV implementation can reach most targets in under 60 steps.
@@
Table 3: VLV VSN experiment. We report the number or successful episodes over 15, the corresponding SR per-object goal, and the average number of actions taken to reach the target.
Figure 10 shows qualitative results for four navigation experiments with the VLV model. We provide five representative images of the navigation experiments. Two successful and two unsuccessful cases are presented. In the first experiment (first row), the robot quickly reached the Table, as the detector easily identified it in the images. In the second experiment, the robot took a detour to the Chair because the detector failed to detect it initially. The model predicted a point near the chair but out of view, causing the robot to move closer only after it became visible. In the third and fourth experiments, the robot started in a challenging position in front of a refrigerator. Due to noise in the depth image, the target point calculation was inaccurate, leading to collisions with the wall in both cases.
@@
We analyze now in detail the results reported by PIRLNav model. Table 4 shows the SR obtained for the PIRLNav model integrated into our ROS4VSN architecture. We detail the SR reported for every object category. The agent was able to more easily locate the most common objects in the house, such as the chair and the monitor. The abundant and well-distributed presence of these objects facilitated the agent’s work. The model inferred the action stop on both objects a total of five times. Large objects, such as the sofa and bed, showed slightly lower results. Although easily visible from multiple locations in the dwelling, the presence of only one of these objects made it difficult for the robotic agent to spot them. The number of times the stop action was sampled for these categories was substantially reduced. With the toilet, we have one of the most complex challenges in navigating the robot in this home. The robot did not have full visibility of the toilet until it managed to fully enter the bathroom, having to pass through the narrow door without colliding. Finally, for the category plant, the robot was not able to locate this category in any of the 15 attempts. Even though the plant was visible on multiple occasions during navigation, the agent did not manage to head towards this object. Overall, considering all the categories, the SR for the PIRLNav model is of $21.11\%$.
@@
Table 4: PIRLNav VSN experiment. We report the number or successful episodes over 15, the corresponding SR per-object goal, and the average number of actions taken to reach the target.